OpenStudy (anonymous):

P= 2L + 2W; L=77 W=80

OpenStudy (anonymous):

and your question is?

OpenStudy (anonymous):

do u want to know what L is ?

OpenStudy (anonymous):

sorry i meant p

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

By plugging in 77 and 80 into the formula

OpenStudy (anonymous):

ok so p = (2*77)+(2*80) P = 154+ 160 = 314

OpenStudy (anonymous):

P = 314

OpenStudy (anonymous):

ok